OpenStudy (anonymous):

**will give medal** What is the speed, in meters per second, of a paper airplane that flies 24 meters in 6 seconds? A. 144 B. 30 C. 18 D. 4

OpenStudy (anonymous):

speed = distance / time s= 24 / 6 s = 4 m/s speed of the airplane is 4m/s

OpenStudy (anonymous):

well because it's per second we wanna find out how many meters for a second so when you set up a proportion, 24/6 = x/1 you should get D as girlbyte said :)